# Non-strict Singularity of Optimal Sobolev Embeddings

### v1

@@ 1. Introduction
where both $X$ and $Y$ are r.i. spaces, $\Omega\subseteq\mathbb{R}^{n}$ is a bounded domain in $\mathbb{R}^{n}$, and $V_{0}^{m}X(\Omega)$ is a homogeneous $m$th-order Sobolev space built upon $X$. They completely characterized the optimal target r.i. space $Y_{X}(\Omega)$ in (1.2). The optimality is meant in the sense that $Y_{X}$ the smallest r.i. space $Y$ with which (1.2) is valid for a given r.i. space $X$. Note that (1.1) is a particular case of (1.2) with $Y(\Omega)=L^{p^{*}}(\Omega)$ and $W^{m,p}_{0}(\Omega)=V_{0}^{m}L^{p}(\Omega)$, up to equivalent norms (see Subsection 2.2 for more information and precise definitions), and the optimal target r.i. space $Y_{X}(\Omega)$ in (1.1) is the Lorentz space $L^{p^{*},p}(\Omega)\subsetneq L^{p^{*}}(\Omega)$ (see O’Neil [1963], Peetre [1966], Tartar [1998]). Furthermore, in the subsequent paper Kerman and Pick [2008], they also completely characterized the compactness of (1.2). In particular, it turns out that the optimal Sobolev embedding $V_{0}^{m}X(\Omega)\hookrightarrow Y_{X}(\Omega)$ is never compact (see also Slavíková [2012]). The interested reader is referred to Breit and Cianchi [2021], Cianchi et al. [2015, 2020] and references therein for more information on optimal Sobolev embeddings in various settings.
@@
In this paper, we develop a general framework for addressing these questions, extending the theory well beyond the $L^{p}$ setting. After recalling the necessary background on r.i. spaces, optimal Sobolev embeddings, and strictly singular operators in Section 2, we introduce in Section 3 the class of sublimiting r.i. norms (Definition 3.1, see also Proposition 3.3). These norms are characterized by the following relation between their fundamental function $\varphi_{X}$ (i.e., the norm of a characteristic function of a set with prescribed measure) and the fundamental function $\varphi_{Y_{X}}$ of the optimal norm in the corresponding Sobolev embedding:
@@
For such norms, we prove a key test-function construction (Theorem 3.6) showing that given an optimal sublimiting Sobolev embedding $V_{0}^{m}X(\Omega)\to Y_{X}(\Omega)$, one can find sequences of radially concentrated “spike” functions whose norms under the Sobolev embedding do not collapse. Notably, this construction bypasses the fact that we do not have any dilation invariance at our disposal for a general optimal embedding $V_{0}^{m}X(\Omega)\to Y_{X}(\Omega)$. Moreover, the optimal target space $Y_{X}$ can be replaced by any r.i. space $Y$ having the same fundamental function as $Y_{X}$.
As our main application (Section 4), we consider the family of Lambda spaces $\Lambda^{q}_{w}$, $q\in[1,\infty)$, defined by
@@
where $w$ is a suitable weight. This class of function spaces encompasses a large number of r.i. spaces, such as Lebesgue spaces, Lorentz spaces, some Orlicz spaces, or Lorentz–Zygmund spaces (see Subsection 2.1). We first show that these norms are sublimiting under natural conditions on the weight $w$ (see Proposition 4.3). Then we show that the corresponding optimal Sobolev embedding
@@
fails to be strictly singular (Theorem 4.5). To achieve this, we leverage our key test-function construction and exploit the known fact that Lambda spaces with exponent $q$ contain an order isomorphic copy of $\ell_{q}$ (see [Kamińska and Maligranda, 2004, Theorem 1]). It should be noted that this known fact still has to be substantially modified and refined to ensure the synergy and coherence of arguments, leading to the desired result. Ultimately, we are able to obtain a rather explicit infinite-dimensional subspace of $V^{m}_{0}\Lambda_{w}^{q}(\Omega)$, isomorphic to $\ell_{q}$, on which the Sobolev embedding (1.3) is an isomorphism onto its image. This shows that the phenomenon observed before only in very special cases of optimal Sobolev embeddings (by using their particular special forms and structures) is prevalent in a large number of general optimal Sobolev embeddings, which can be described as (1.3).
@@ 2.1. Rearrangement-invariant function spaces
We always have $f^{*}\leq f^{**}$. Unlike $f^{*}$, $f^{**}$ need not be equimeasurable with $f$.
@@
The rearrangement-invariant function space (r.i. space for short) $X(R,\mu)$ is defined as
@@
A function $\varphi\colon[0,1]\to[0,\infty)$ is quasiconcave if $\varphi$ is nondecreasing, the function $(0,1]\ni t\mapsto\varphi(t)/t$ is nonincreasing, and $\varphi(t)=0$ if and only if $t=0$. A quasiconcave function is always continuous on $(0,1]$, but it may have a jump at 00. For example, if $X(R,\mu)=L^{p}(R,\mu)$, then $\varphi_{X}(t)=t^{1/p}$, $t\in(0,1]$. The fundamental function of an r.i. space $X(R,\mu)$ need not be concave, but $X(R,\mu)$ can always be equivalently renormed with another r.i. function norm in such a way that the resulting fundamental function is concave. More specifically, we can do it in such a way that the resulting fundamental function is the least nondecreasing concave majorant of $\varphi_{X}$. Given a quasiconcave function $\varphi\colon[0,1]\to[0,\infty)$, its least nondecreasing concave majorant $\tilde{\varphi}$ satisfies
@@
[…]
@@
See Bennett and Rudnick [1980], Opic and Pick [1999] for more information. Recall that a function $A\colon[0,\infty]\to[0,\infty]$ is called a Young function if it is convex, left-continuous, vanishing at 00, and not constant on the entire interval $(0,\infty)$. Given a Young function $A$, the Orlicz space $L^{A}(R,\mu)$ is defined as the collection of those $f\in\mathscr{M}(R,\mu)$ for which
@@
Let $X(\Omega)$ be an r.i. space and $m\in\mathbb{N}$. The Sobolev space $V^{m}_{0}X(\Omega)$ is defined as the space of all $m$-times weakly differentiable functions $u$ on $\Omega$ whose continuation by 00 outside $\Omega$ is an $m$-times weakly differentiable function and
@@ 3. Optimal Sublimiting Embeddings
The main result of this section is Theorem 3.6, which will play an important role in the proof of our main result, Theorem 4.5. To prove Theorem 3.6, we first need to establish a few results regarding the optimal Sobolev embedding (2.15), which are of independent interest. We start with a proposition telling us that the “small” and “capital” endpoint spaces $m$ and $M$ corresponding to optimal spaces in Sobolev embeddings (2.15) coincide.
@@
    Since the functions on both sides of (3.7) are positive and continuous on the interval $[\delta,1]$ for each $\delta\in(0,1)$, in order to establish the equivalence, it is sufficient to establish it for every $t\in(0,\delta)$ for some $\delta>0$.
@@
There are borderline cases that are not sublimiting in the sense of Definition 3.1 and at the same time, nor are they (super)limiting in the sense that $V^{m}_{0}X(\Omega)$ is embedded in $L^{\infty}(\Omega)$. In particular, the latter means that (3.9) is satisfied for such spaces. Typical examples of such borderline cases are Lorentz spaces $X=L^{\frac{n}{m},q}$ with $q\in(1,\infty]$. Then $Y_{X}(\Omega)$ is the Lorentz–Zygmund space $L^{\infty,q,-1}(\Omega)$ (e.g., see [Cianchi and Pick, 2016, Theorem 5.1]), whose fundamental function satisfies
@@
Our next proposition is a technical result, which will come handy in the proof of Theorem 3.6. We first need to make some definitions and recall an important result. For $a\in(0,1]$, we define
@@
(3.25) the support of uf is the closed ball $\text{the support of $u_{f}$ is the closed ball $\bar{B}(x_{0},a^{\frac{1}{n}}R)$}.$
@@
thanks to Proposition 3.5. By [Kerman and Pick, 2006, Theorem A] combined with (3.22), we have
@@
Next, let $B(x_{0},R)\subseteq\Omega$ be a ball inside $\Omega$. By following and suitably modifying the construction from [Kerman and Pick, 2006, p. 562] (cf. [Mihula, , Proposition 4.5]), we can construct for every $f\in L^{\infty}(0,1)$ that has support in $[0,\delta]$ for some $\delta\in(0,1)$ functions $u_{f}$ and $u_{|f|}$ such that both $u_{f}$ and $u_{|f|}$ belong to $V^{m}_{0}X(\Omega)\cap\mathcal{C}^{m-1}(\Omega\setminus\{x_{0}\})$, $u_{|f|}$ is radially nonincreasing with respect to $x_{0}$, $|u_{f}|\leq u_{|f|}$,
@@
(3.32) the support of u|f| is the closed ball $\text{the support of $u_{|f|}$ is the closed ball $\bar{B}(x_{0},\varrho^{\frac{1}{n}}R)$},$
@@
and in which the constants $C_{1}$ and $C$ depend only on $m$, $n$, and $R$ (in particular, they are independent of $\delta$ and $f$). The fact that $u_{|f|}\in Y(\Omega)$ follows from (3.22). Moreover, the assignment $f\mapsto u_{f}$ is linear, i.e., if $\sum_{k=1}^{M}\alpha_{j}f_{j}$ is a linear combination of functions from $L^{\infty}(0,1)$ having supports in $[0,\delta]$ for some $\delta\in(0,1)$, then
@@
$\nabla u_{f}(x)=\frac{n}{R}f\Big{(}\frac{|x-x_{0}|^{n}}{R^{n}}\Big{)}\frac{x-x_{0}}{|x-x_{0}|}\chi_{B(x_{0},R)}(x)\quad\text{for a.e.\leavevmode\nobreak\ $x\in\Omega$}.$
@@ 4. Optimal embeddings of “nonlimiting” $\Lambda$ spaces and their non-strict singularity
We start with an auxiliary proposition of independent interest, which will be an essential ingredient in the proof of our main result, Theorem 4.5. Loosely speaking, it tells us that when $X$ is a Lambda space, the optimal r.i. space in Sobolev embeddings (2.15) that are not (super)limiting is a Lambda space of the same type. The price for the generality of Proposition 4.1 is that its assumptions are unavoidably technical. Nevertheless, it is usually not difficult to verify them (see Remark 4.2).
@@
First, let $q=1$. Since the weight $w$ satisfies (2.12), we may without loss of generality assume that the function $W$ is concave. Hence, it follows from (2.11) that
@@
$w_{opt}(t)=\Bigg{(}\Big{(}1+\int_{s}^{1}\frac{w(\tau)}{W(\tau)^{q^{\prime}}}\tau^{\frac{m}{n}q^{\prime}}\mathop{}\!\mathrm{d}{\tau}\Big{)}^{1-q}\Bigg{)}^{\prime}(t)\quad\text{for a.e.\leavevmode\nobreak\ $t\in(0,1)$},$
@@
    As explained at the beginning of the proof of Proposition 4.1, the assumptions (4.1) and (4.2) are equivalent to the fact that $X=\Lambda^{q}_{w}$ is not (super)limiting, by which we mean that $V^{m}_{0}X(\Omega)$ is not embedded in $L^{\infty}(\Omega)$. When $q\in(1,\infty)$, the assumptions (4.3) and (4.4) ensure, loosely speaking, that $X=\Lambda^{q}_{w}$ is not “too close to $L^{1}$” and not “too close to $L^{\infty}$”, respectively. Typically, the validity of (4.4) is already implied by that of (4.2).
@@
    Consider the Lorentz spaces $L^{p,q}(0,1)=\Lambda^{q}_{w_{p,q}}(0,1)$ (recall (2.14)). Note that in order that $w$ is a weight, we need to have $p<\infty$. Moreover, when $q\in(1,\infty)$, (4.4) is satisfied for $p<\infty$. It is easy to see that (2.12) is satisfied if and only if either $p\in(1,\infty)$ and $q\in[1,\infty)$ or $p=q=1$. When $q=1$, the condition (4.1) is satisfied if and only if $p<n/m$. When $q\in(1,\infty)$, the condition (4.2) is satisfied if and only if $p\leq n/m$, whereas (4.3) is satisfied if and only if $p>1$. Putting everything together, we see that the assumptions of Proposition 4.1 are satisfied if and only if either $q=1$ and $p\in[1,n/m)$ or $q\in(1,\infty)$ and $p\in(1,n/m]$.
@@
    The assumptions of Proposition 4.1 do not guarantee that $\|\cdot\|_{X(0,1)}=\|\cdot\|_{\Lambda^{q}_{w}(0,1)}$ is sublimiting in the sense of Definition 3.1. For example, they are satisfied for the Lorentz spaces $L^{\frac{n}{m},q}(0,1)=\Lambda^{q}_{w_{n/m,q}}(0,1)$ with $q\in(1,\infty)$ (see the preceding remark and Remark 3.4).
The following proposition is essentially a particular case of Proposition 4.1 with the extra information that $\|\cdot\|_{X(0,1)}=\|\cdot\|_{\Lambda^{q}_{w}(0,1)}$ is sublimiting in the sense of Definition 3.1.
@@
In view of Proposition 4.1, Remark 3.2(ii), and (3.7), we only need to show that the validity of (4.21) implies the validity of (4.1) or (4.2). Suppose that this is not the case. As explained at the beginning of the proof of Proposition 4.1, this means that the embedding
@@
Furthermore, as in the proof of Proposition 3.3, the validity of (4.23) is equivalent to the fact that $Y_{X}=L^{\infty}$. It follows that
@@
Combining this, the fact that $\|\cdot\|_{X(0,1)}=\|\cdot\|_{\Lambda^{q}_{w}(0,1)}$ is sublimiting (recall Remark 3.2(ii)), and (3.7), we obtain
@@
We are finally ready to prove our main result, which loosely speaking tells us that the optimal sublimiting Sobolev embeddings (2.15) for Lambda spaces are not strictly singular. Its proof rests on Theorem 3.6 and Proposition 4.3, which are used in such a way that we can exploit the fact that both $X=\Lambda^{q}_{w}$ and $Y_{X}=\Lambda^{q}_{w_{opt}}$ contain an order isomorphic copy of $\ell_{q}$ (see [Kamińska and Maligranda, 2004, Theorem 1]).
@@
By Proposition 4.3, $\|\cdot\|_{X(0,1)}=\|\cdot\|_{\Lambda^{q}_{w}(0,1)}$ is sublimiting and the optimal target r.i. space $Y_{X}(\Omega)$ satisfies (4.5) with a weight $w_{opt}$ for which (4.22) is true. Furthermore, it is easy to see that the fact that the optimal embedding (4.26) is not strictly singular follows from the existence of a linear independent sequence of functions $\{u_{j}\}_{j=1}^{\infty}\subseteq V^{m}_{0}\Lambda_{w}^{q}(\Omega)$ for which (4.27) and (4.28) are true (with an arbitrarily chosen $\varepsilon_{1},\varepsilon_{2}>0$ in such a way that $C_{2}/(1+\varepsilon_{1})-\varepsilon_{2}>0$). Therefore, we only need to prove the existence of such a sequence of functions.
@@
Fix a ball $B=B(x_{0},R)\subseteq\Omega$ inside $\Omega$. Let $\varepsilon_{1},\varepsilon_{2}>0$ be given. We will construct the desired linear independent sequence of functions $\{u_{j}\}_{j=1}^{\infty}\subseteq V^{m}_{0}\Lambda_{w}^{q}(\Omega)$ inductively. Set $a_{1}=1/2$ and $B_{1}=B(x_{0},a_{1}^{1/n}R)$. By Theorem 3.6, there are positive constants $C_{1}$ and $C_{2}$, and functions $f_{1}\in\mathscr{S}_{a_{1}/2,a_{1}}$ and $u_{1}=u_{f_{1}}\in V^{m}_{0}X(\Omega)\cap\mathcal{C}^{m-1}(\Omega\setminus\{x_{0}\})$ such that
@@
(4.35) the support of u1 is the closed ball $\text{the support of $u_{1}$ is the closed ball $\bar{B}_{1}$}.$

### v2

@@ 1. Introduction
where both $X$ and $Y$ are r.i. spaces, $\Omega\subseteq\mathbb{R}^{n}$ is a bounded domain in $\mathbb{R}^{n}$, and $V_{0}^{m}X(\Omega)$ is a homogeneous $m$th-order Sobolev space built upon $X$. They completely characterized the optimal target r.i. space $Y(\Omega)$ in (1.2). We will denote this optimal space by $Y_{X}$. The optimality is meant in the sense that $Y_{X}$ the smallest r.i. space $Y$ with which (1.2) is valid for a given r.i. space $X$. Note that (1.1) is a particular case of (1.2) with $Y(\Omega)=L^{p^{*}}(\Omega)$ and $W^{m,p}_{0}(\Omega)=V_{0}^{m}L^{p}(\Omega)$, up to equivalent norms (see Subsection 2.2 for more information and precise definitions), and the optimal target r.i. space $Y_{X}(\Omega)$ in (1.1) is the Lorentz space $L^{p^{*},p}(\Omega)\subsetneq L^{p^{*}}(\Omega)$ (see O’Neil [1963], Peetre [1966], Tartar [1998]). Furthermore, in the subsequent paper Kerman and Pick [2008], they also completely characterized the compactness of (1.2). In particular, it turns out that the optimal Sobolev embedding $V_{0}^{m}X(\Omega)\hookrightarrow Y_{X}(\Omega)$ is never compact (see also Cavaliere and Mihula [2019, 2022], Slavíková [2012, 2015]). The interested reader is referred to Breit and Cianchi [2021], Cianchi et al. [2015, 2020] and references therein for more information on optimal Sobolev embeddings in various settings.
@@
In this paper, we develop a general framework for addressing these questions, extending the theory well beyond the $L^{p}$ setting. After recalling the necessary background on r.i. spaces, optimal Sobolev embeddings, and strictly singular operators in Section 2, we introduce in Section 3 the class of sublimiting r.i. norms (Definition˜3.2, see also Proposition˜3.4). These norms are characterized by the following relation between their fundamental function $\varphi_{X}$ (i.e., the norm of a characteristic function of a set with prescribed measure) and the fundamental function $\varphi_{Y_{X}}$ of the optimal norm in the corresponding Sobolev embedding:
@@
For such norms, we prove a key test-function construction (Theorem˜3.7) showing that given an optimal sublimiting Sobolev embedding $V_{0}^{m}X(\Omega)\to Y_{X}(\Omega)$, one can find sequences of radially concentrated “spike” functions whose norms under the Sobolev embedding do not collapse. Notably, this construction bypasses the fact that we do not have any dilation invariance at our disposal for a general optimal embedding $V_{0}^{m}X(\Omega)\to Y_{X}(\Omega)$. Moreover, the optimal target space $Y_{X}$ can be replaced by any r.i. space $Y$ having the same fundamental function as $Y_{X}$.
As our main application (Section 4), we consider the family of classical Lorentz spaces of type Lambda (Lambda spaces, for short) $\Lambda^{q}_{w}$, $q\in[1,\infty)$, defined by
@@
where $w$ is a suitable weight. This class of function spaces encompasses a large number of r.i. spaces, such as Lebesgue spaces, Lorentz spaces, some Orlicz spaces, or Lorentz–Zygmund spaces (see Subsection 2.1). We first show that these norms are sublimiting under natural conditions on the weight $w$ (see Proposition˜4.3). Then we show that the corresponding optimal Sobolev embedding
@@
fails to be strictly singular (Theorem˜4.5), where $w_{opt}$ is an optimal weight (see the proof of Proposition˜4.1 for its definition). To achieve this, we leverage our key test-function construction and exploit the known fact that Lambda spaces with exponent $q$ contain an order isomorphic copy of $\ell_{q}$ (see [Kamińska and Maligranda, 2004, Theorem 1]). It should be noted that this known fact still has to be substantially modified and refined to ensure the synergy and coherence of arguments, leading to the desired result. Ultimately, we are able to obtain a rather explicit infinite-dimensional subspace of $V^{m}_{0}\Lambda_{w}^{q}(\Omega)$, isomorphic to $\ell_{q}$, on which the Sobolev embedding (1.3) is an isomorphism onto its image. This shows that the phenomenon observed before only in very special cases of optimal Sobolev embeddings (by using their particular special forms and structures) is prevalent in a large number of general optimal Sobolev embeddings, which can be described as (1.3).
@@ 2.1. Rearrangement-invariant function spaces
We always have $f^{*}\leq f^{**}$. Unlike $^{*}, ff^{**}$ need not be equimeasurable with $f$.
@@
The rearrangement-invariant space (r.i. space for short) $X(R,\mu)$ is defined as
@@
A function $\varphi\colon[0,1]\to[0,\infty)$ is quasiconcave if $\varphi$ is nondecreasing, the function $(0,1]\ni t\mapsto\varphi(t)/t$ is nonincreasing, and $\varphi(t)=0$ if and only if $t=0$. A quasiconcave function is always continuous on $(0,1]$, but it may have a jump at 0. For example, if $X(R,\mu)=L^{p}(R,\mu)$, then $\varphi_{X}(t)=t^{1/p}$, $t\in(0,1]$. The fundamental function of an r.i. space $X(R,\mu)$ need not be concave, but $X(R,\mu)$ can always be equivalently renormed with another r.i. function norm in such a way that the resulting fundamental function is concave. More specifically, we can do it in such a way that the resulting fundamental function is the least nondecreasing concave majorant of $\varphi_{X}$. Given a quasiconcave function $\varphi\colon[0,1]\to[0,\infty)$, its least nondecreasing concave majorant $\tilde{\varphi}$ satisfies
@@
if and only if (e.g., see [Musil and Oľhava, 2019, Lemma 2.1] and references therein)
@@
See Bennett and Rudnick [1980], Opic and Pick [1999] for more information. Recall that a function $A\colon[0,\infty]\to[0,\infty]$ is called a Young function if it is convex, left-continuous, vanishing at 0, and not constant on the entire interval $(0,\infty)$. Given a Young function $A$, the Orlicz space $L^{A}(R,\mu)$ is defined as the collection of those $f\in\mathscr{M}(R,\mu)$ for which
@@
Let $X(\Omega)$ be an r.i. space and $m\in\mathbb{N}$. The Sobolev space $V^{m}_{0}X(\Omega)$ is defined as the space of all $m$-times weakly differentiable functions $u$ on $\Omega$ whose continuation by 0 outside $\Omega$ is an $m$-times weakly differentiable function and
@@ 3. Optimal Sublimiting Embeddings
The main result of this section is Theorem˜3.7, which will play an important role in the proof of our main result, Theorem˜4.5. To prove Theorem˜3.7, we first need to establish a few results regarding the optimal Sobolev embedding (2.15), which are of independent interest. We start with a proposition telling us that the “small” and “capital” endpoint spaces $m$ and $M$ corresponding to optimal spaces in Sobolev embeddings (2.15) coincide.
@@
    Since the functions on both sides of (3.7) are positive and continuous on the interval $\delta,1]$ for each $\delta\in(0,1)$, in order to establish the equivalence, it is sufficient to establish it for every $t\in(0,\delta)$ for some $\delta>0$.
@@
There are borderline cases that are not sublimiting in the sense of Definition˜3.2 and at the same time, nor are they (super)limiting in the sense that $V^{m}_{0}X(\Omega)$ is embedded in $L^{\infty}(\Omega)$. In particular, the latter means that (3.9) is satisfied for such spaces. Typical examples of such borderline cases are Lorentz spaces $X=L^{\frac{n}{m},q}$ with $q\in(1,\infty]$. Then $Y_{X}(\Omega)$ is the Lorentz–Zygmund space $L^{\infty,q,-1}(\Omega)$ (e.g., see [Cianchi and Pick, 2016, Theorem 5.1]), whose fundamental function satisfies
@@
Our next proposition is a technical result, which will come handy in the proof of Theorem˜3.7. We first need to make some definitions and recall an important result. For $a\in(0,1]$, we define
@@
[…]
@@
thanks to Proposition˜3.6. By [Kerman and Pick, 2006, Theorem A] combined with (3.22), we have
@@
Next, let $B(x_{0},R)\subseteq\Omega$ be a ball inside $\Omega$. By following and suitably modifying the construction from [Kerman and Pick, 2006, p. 562] (cf. [Mihula, , Proposition 4.5]), we can construct for every $f\in L^{\infty}(0,1)$ that has support in $0,\delta]$ for some $\delta\in(0,1)$ functions $u_{f}$ and $u_{|f|}$ such that both $u_{f}$ and $u_{|f|}$ belong to $V^{m}_{0}X(\Omega)\cap\mathcal{C}^{m-1}(\Omega\setminus\{x_{0}\})$, $u_{|f|}$ is radially nonincreasing with respect to $x_{0}$, $|u_{f}|\leq u_{|f|}$,
@@
[…]
@@
and in which the constants $C_{1}$ and $C$ depend only on $m$, $n$, and $R$ (in particular, they are independent of $\delta$ and $f$). The fact that $u_{|f|}\in Y(\Omega)$ follows from (3.22). Moreover, the assignment $f\mapsto u_{f}$ is linear, i.e., if $\sum_{k=1}^{M}\alpha_{j}f_{j}$ is a linear combination of functions from $L^{\infty}(0,1)$ having supports in $0,\delta]$ for some $\delta\in(0,1)$, then
@@
$\nabla u_{f}(x)=\frac{n}{R}f\Big(\frac{|x-x_{0}|^{n}}{R^{n}}\Big)\frac{x-x_{0}}{|x-x_{0}|}\chi_{B(x_{0},R)}(x)\quad\text{for a.e.\penalty 10000\ $x\in\Omega$}.$
@@ 4. Optimal embeddings of “nonlimiting” $\Lambda$ spaces and their non-strict singularity
We start with an auxiliary proposition of independent interest, which will be an essential ingredient in the proof of our main result, Theorem˜4.5. Loosely speaking, it tells us that when $X$ is a Lambda space, the optimal r.i. space in Sobolev embeddings (2.15) that are not (super)limiting is a Lambda space of the same type. The price for the generality of Proposition˜4.1 is that its assumptions are unavoidably technical. Nevertheless, it is usually not difficult to verify them (see Remark˜4.2).
@@
First, let $q=1$. Since the weight $w$ satisfies (2.12), we may assume without loss of generality that the function $W$ is concave. Hence, it follows from (2.11) that
@@
$w_{opt}(t)=\Bigg(\Big(1+\int_{s}^{1}\frac{w(\tau)}{W(\tau)^{q^{\prime}}}\tau^{\frac{m}{n}q^{\prime}}\mathop{}\!\mathrm{d}{\tau}\Big)^{1-q}\Bigg)^{\prime}(t)\quad\text{for a.e.\penalty 10000\ $t\in(0,1)$},$
@@
    As explained at the beginning of the proof of Proposition˜4.1, the assumptions (4.1) and (4.2) are equivalent to the fact that $X=\Lambda^{q}_{w}$ is not (super)limiting, by which we mean that $V^{m}_{0}X(\Omega)$ is not embedded in $L^{\infty}(\Omega)$. When $q\in(1,\infty)$, the assumptions (4.3) and (4.4) ensure, loosely speaking, that $X=\Lambda^{q}_{w}$ is not “too close to $L^{1}$” and not “too close to $L^{\infty}$”, respectively. Typically, the validity of (4.4) is already implied by that of (4.2).
@@
    Consider the Lorentz spaces $L^{p,q}(0,1)=\Lambda^{q}_{w_{p,q}}(0,1)$ (recall (2.14)). Note that in order that $w$ is a weight, we need to have $p<\infty$. Moreover, when $q\in(1,\infty)$, (4.4) is satisfied for $p<\infty$. It is easy to see that (2.12) is satisfied if and only if either $p\in(1,\infty)$ and $q\in[1,\infty)$ or $p=q=1$. When $q=1$, the condition (4.1) is satisfied if and only if $p<n/m$. When $q\in(1,\infty)$, the condition (4.2) is satisfied if and only if $p\leq n/m$, whereas (4.3) is satisfied if and only if $p>1$. Putting everything together, we see that the assumptions of Proposition˜4.1 are satisfied if and only if either $q=1$ and $p\in[1,n/m)$ or $q\in(1,\infty)$ and $p\in(1,n/m]$.
@@
    The assumptions of Proposition˜4.1 do not guarantee that $\|\cdot\|_{X(0,1)}=\|\cdot\|_{\Lambda^{q}_{w}(0,1)}$ is sublimiting in the sense of Definition˜3.2. For example, they are satisfied for the Lorentz spaces $L^{\frac{n}{m},q}(0,1)=\Lambda^{q}_{w_{n/m,q}}(0,1)$ with $q\in(1,\infty)$ (see the preceding remark and Remark˜3.5).
The following proposition is essentially a particular case of Proposition˜4.1 with the extra information that $\|\cdot\|_{X(0,1)}=\|\cdot\|_{\Lambda^{q}_{w}(0,1)}$ is sublimiting in the sense of Definition˜3.2.
@@
In view of Proposition˜4.1, Remark˜3.3(ii), and (3.7), we only need to show that the validity of (4.21) implies that of (4.1) or (4.2). Suppose that this is not the case. As explained at the beginning of the proof of Proposition˜4.1, this means that the embedding
@@
Furthermore, as in the proof of Proposition˜3.4, the validity of (4.23) is equivalent to the fact that $Y_{X}=L^{\infty}$. It follows that
@@
Combining this, the fact that $\|\cdot\|_{X(0,1)}=\|\cdot\|_{\Lambda^{q}_{w}(0,1)}$ is sublimiting (recall Remark˜3.3(ii)), and (3.7), we obtain
@@
We are finally ready to prove our main result, which loosely speaking tells us that the optimal sublimiting Sobolev embeddings (2.15) for Lambda spaces are not strictly singular. Its proof rests on Theorem˜3.7 and Proposition˜4.3, which are used in such a way that we can exploit the fact that both $X=\Lambda^{q}_{w}$ and $Y_{X}=\Lambda^{q}_{w_{opt}}$ contain an order isomorphic copy of $\ell_{q}$ (see [Kamińska and Maligranda, 2004, Theorem 1]).
@@
By Proposition˜4.3, $\|\cdot\|_{X(0,1)}=\|\cdot\|_{\Lambda^{q}_{w}(0,1)}$ is sublimiting and the optimal target r.i. space $Y_{X}(\Omega)$ satisfies (4.5) with a weight $w_{opt}$ for which (4.22) is true. Furthermore, it is easy to see that the fact that the optimal embedding (4.26) is not strictly singular follows from the existence of a linear independent sequence of functions $\{u_{j}\}_{j=1}^{\infty}\subseteq V^{m}_{0}\Lambda_{w}^{q}(\Omega)$ for which (4.27) and (4.28) are true (with an arbitrarily chosen $\varepsilon_{1},\varepsilon_{2}>0$ in such a way that $C_{2}/(1+\varepsilon_{1})-\varepsilon_{2}>0$). Therefore, we only need to prove the existence of such a sequence of functions.
@@
Fix a ball $B=B(x_{0},R)\subseteq\Omega$ inside $\Omega$. Let $\varepsilon_{1},\varepsilon_{2}>0$ be given. We will construct the desired linear independent sequence of functions $\{u_{j}\}_{j=1}^{\infty}\subseteq V^{m}_{0}\Lambda_{w}^{q}(\Omega)$ inductively. Set $a_{1}=1/2$ and $B_{1}=B(x_{0},a_{1}^{1/n}R)$. By Theorem˜3.7, there are positive constants $C_{1}$ and $C_{2}$, and functions $f_{1}\in\mathscr{S}_{a_{1}/2,a_{1}}$ and $u_{1}=u_{f_{1}}\in V^{m}_{0}X(\Omega)\cap\mathcal{C}^{m-1}(\Omega\setminus\{x_{0}\})$ such that
@@
[…]